Assume that an event A contains 10 observations and event B contains 15 observations. If the intersection of events A and B contains exactly 3 observations, how many observations are in the union of these two events? A. 0 B. 10 C. 22 D. 28
Found 2 solutions by sudhanshu_kmr, ewatrrr:
Answer by sudhanshu_kmr(1152)
Answer by ewatrrr(24785)

 
Hi,
event A contains 10 observations and event B contains 15 observations.
intersection of events A and B contains exactly 3
how many observations are in the union of these two events:
P(A ∪ B) = P(A) + P(B) - P(A ∩ B) = 10 + 15 - 3 = 22